The sum of two numbers is 70. The second is 10 more than 3 times the first. Find the numbers. Choose the answer with the correct equations used to solve this problem.
Answer by stanbon(48568)

The sum of two numbers is 70. The second is 10 more than 3 times the first. Find the numbers. Choose the answer with the correct equations used to solve this problem.
----------------
Equations:
x + y = 70
y = 3x+10
------------------
Subsitute for "y" and solve for "x":
x + (3x+10) = 70
4x = 60
x = 15 (1st number)
y = 3x+10, so y = 45*10 = 55 (2nd number)
==========================================
Cheers,
Stan H.